OpenStudy (anonymous):

olve: m/(m-7)=(m/(m-7)) + 11

OpenStudy (anonymous):

There are no solutions.

OpenStudy (anonymous):

Thanks that what I figured but I was just checking

OpenStudy (anonymous):

Can't be. It's like x=x+11, with x = m/(m-7)